# Policy Learning with Confidence

### v2

@@ Abstract.
This paper introduces a framework for selecting policies that maximize expected welfare under estimation uncertainty. The proposed method explicitly balances the size of the estimated welfare against the uncertainty inherent in its estimation, ensuring that chosen policies meet a reporting guarantee, namely, that actual welfare is guaranteed not to fall below the reported estimate with a pre-specified confidence level. We produce the efficient decision frontier, describing policies that offer maximum estimated welfare for a given acceptable level of estimation risk. We apply this approach to a variety of settings, including the selection of policy rules that allocate individuals to treatments and the allocation of limited budgets among competing social programs.
We thank Isaiah Andrews, Nathan Hendren, Matt Masten, Whitney Newey, Vira Semenova, Jesse Shapiro, Aleksey Tetenov and conference and seminar participants at the Munich Econometrics Workshop, Duke University, CUNY, University of Geneva, Tilburg University, the Encounters in Econometric Theory workshop, CUHK Shenzhen, and IAAE for helpful comments and discussion. Tian Xie provided excellent research assistance. Liyang Sun acknowledges generous support from the Economic and Social Research Council (new investigator grant UKRI607).
@@ 1. Introduction
Consider a decision-maker (DM) faced with choosing from a menu of policies to maximize expected welfare. The DM could for example be a planner deciding how to allocate treatments to heterogeneous individuals, a firm choosing different potential innovations in which to invest resources, an auctioneer choosing an optimal auction design or reserve price, or a policy-maker choosing expenditure shares to allocate to government programs to maximize a measure of welfare to the public. If the DM knew the welfare that would be obtained from implementing each policy, the choice would be easy; she would then simply choose the policy that would yield the highest welfare. However, the DM lacks such precise knowledge, but instead has data that can be used to consistently estimate the welfare that would be achieved by each policy. The estimates are measured with varying degrees of precision, as reflected by their standard errors. How should the DM use the available information for policy selection?
An intuitively appealing choice is to select the policy with the highest estimated welfare, the so-called plug-in rule, which simply replaces the population objective with its sample analog in determining policy choice. In the literature on statistical treatment rules, this is referred to as empirical welfare maximization (EWM). [^1] See for example Kitagawa and Tetenov (2018) and Athey and Wager (2021). Andrews et al. (2024) alternatively refers to this as the “natural rule” or “picking the winners”. The same logical task applies more broadly to a variety of economic contexts, for example when researchers use structural models, they obtain estimates of the performance of different policies that can then be used to inform policy choice. Manski (2021) calls this “as-if” optimization, which he describes as, “specification of a model, point estimation of its parameters, and use of the point estimate to make a decision that would be optimal if the estimate were accurate.” Here we refer to such rules interchangeably as plug-in rules or empirical welfare maximization (EWM) rules.
@@
This paper proposes a framework for selecting policies with the goal of maximizing expected welfare in the presence of estimation uncertainty, explicitly accounting for estimation risk in the decision rule. As inputs to her decision the DM has consistent estimates for the welfare of each available policy, and their corresponding sample variance matrix, with sample estimates approximately normally distributed around their population means. [^2] If the estimates are from independent samples the sample variance matrix will simply be a diagonal matrix with each policy’s sample variance along the diagonal. We define and analyze a class of risk-aware policy rules that provide a principled manner for explicitly balancing the estimated welfare of each policy against its estimation risk. Policies that perform well in sample but are estimated imprecisely may be passed over in favor of less risky policies that are not optimal according to the EWM rule but that can be more confidently assured to achieve high performance. Risk-aware rules select a policy from the (Pareto) efficient decision frontier that balances performance and precision as measured by trading off higher sample estimates against smaller sample variance, where the exact tradeoff is governed by the DM’s choice of policy rule from the class of risk-aware rules. Consequently, such rules select the best performing policy among those within the DM’s choice of acceptable estimation error. They can alternatively be interpreted as selection rules obtained by a risk-averse Bayesian DM who minimizes an expected loss function that takes a weighted average of regret and estimation error. [^3] This is shown in Section 2.5. Such dual-objective loss functions have been explored in the statistics literature in e.g. Gupta and Miescke (1986), Gupta and Miescke (1990), and see also Liese and Miescke (2008), as a way to unify the problems of selection and estimation when sampling from independent normal populations. We establish favorable regret properties for this class of rules, as well as conditions under which they deliver a reporting guarantee, ensuring with high confidence that the actual welfare exceeds a lower threshold.
@@
To illustrate, consider a setting in which a budget-conscious DM needs to allocate funds across several social programs that affect the same group of individuals, either in response to a slight budget increase or for incremental cost-cutting. Suppose the DM wishes to use the Marginal Value of Public Funds (MVPF), spear-headed by Hendren and Sprung-Keyser (2020), for this purpose. The MVPF of each social program is the ratio of its marginal benefit to its net marginal cost to the government, inclusive of the impact of any behavioral responses on the government budget. For instance, an MVPF of 1.5 indicates that each \$1 of net government spending generates \$1.50 in benefits for beneficiaries. For a utilitarian DM, the MVPF provides a metric to compare the “bang for the buck” of different policies that target the same beneficiaries. A DM interested in allocating additional marginal expenditure will prioritize programs with the highest MVPF to maximize welfare. [^4] Here we abstract from distributional incidence and assume the DM values dollars equally across the beneficiaries of each policy. If welfare weights vary across beneficiaries, as in the original framework of Hendren and Sprung-Keyser (2020), our analysis can accommodate this by adjusting each MVPF in the value function using the DM’s designated weights. Example welfare weights have been derived in Hendren (2020). Other metrics for measuring program’s marginal benefit and cost have also been discussed in García and Heckman (2022).

### v3

@@ Abstract.
This paper introduces a rule for policy selection in the presence of estimation uncertainty, explicitly accounting for estimation risk. The rule belongs to the class of risk-aware rules on the efficient decision frontier, characterized as policies offering maximal estimated welfare for a given level of estimation risk. Among this class, the proposed rule is chosen to provide a reporting guarantee, ensuring that the welfare delivered exceeds a threshold with a pre-specified confidence level. We apply this approach to the allocation of a limited budget among social programs using estimates of their marginal value of public funds and associated standard errors.
This is a revised version of the arXiv working paper 2502.10653 originally posted in February 2025. We thank Isaiah Andrews, Jiafeng (Kevin) Chen, Nathan Hendren, Toru Kitagawa, Matt Masten, Adam McCloskey, Whitney Newey, Vira Semenova, Jesse Shapiro, Aleksey Tetenov and conference and seminar participants at the Munich Econometrics Workshop, Duke University, CUNY, University of Geneva, Tilburg University, the Encounters in Econometric Theory workshop, CUHK Shenzhen, IAAE, Goethe University Frankfurt, the 2025 ESWC, and the 2026 ASSA meeting for helpful comments and discussion. Tian Xie provided excellent research assistance. Liyang Sun acknowledges generous support from the Economic and Social Research Council (new investigator grant UKRI607).
@@ 1. Introduction
Consider a decision maker (DM) faced with choosing from a menu of policies to maximize expected welfare. The DM could be a planner deciding how to allocate treatments to heterogeneous individuals, a firm choosing different potential innovations in which to invest resources, an auctioneer choosing an optimal auction design or reserve price, or a policy-maker choosing expenditure shares to allocate to government programs to maximize a measure of public welfare. If the DM knew the welfare that would be obtained from implementing each policy, the choice would be easy; she would then simply choose the policy that would yield the highest welfare. However, the DM lacks such precise knowledge, but instead has data from one or possibly multiple studies that can be used to consistently estimate the welfare that would be achieved by each policy. The estimates are measured with varying degrees of precision, as reflected by their standard errors. How should the DM use the available information for policy selection?
An intuitively appealing choice is to select the policy with the highest estimated welfare, the so-called plug-in rule, which simply replaces the population objective with its sample analog in determining policy choice. In the literature on policy learning, this is referred to as empirical welfare maximization (EWM). [^1] See for example Kitagawa and Tetenov (2018) and Athey and Wager (2021). Andrews et al. (2024) alternatively refers to this as the “natural rule” or “picking the winners”. The same logical task applies more broadly to a variety of economic contexts, for example when researchers use structural models, they obtain estimates of the performance of different policies that can then be used to inform policy choice. Manski (2021) calls this “as-if” optimization, described as, “specification of a model, point estimation of its parameters, and use of the point estimate to make a decision that would be optimal if the estimate were accurate.” Here we refer to such rules interchangeably as plug-in rules or empirical welfare maximization (EWM) rules.
@@
This paper proposes a rule for selecting policies with the goal of maximizing expected welfare in the presence of estimation uncertainty, explicitly accounting for estimation risk in the decision rule. As inputs to her decision the DM has consistent estimates for the welfare of each available policy, and their corresponding sample variance matrix, with sample estimates approximately normally distributed around their population means. [^2] If the estimates are from independent samples the sample variance matrix will simply be a diagonal matrix with each policy’s sample variance along the diagonal. We define and analyze a class of risk-aware policy rules that provide a principled manner for explicitly balancing the estimated welfare of each policy against its estimation risk, and which we show have favorable regret properties. Risk-aware rules select a policy from the (Pareto) efficient decision frontier that balances performance and precision as measured by trading off higher sample estimates against smaller sample variance, where the exact tradeoff is governed by the DM’s choice of policy rule from the class of risk-aware rules. Our newly proposed rule is determined from within this class by balancing this tradeoff using the tangent line to the frontier whose slope is determined by the critical value of a one-sided upper confidence band for welfare following the approach for intersection bound inference developed in Chernozhukov et al. (2013). Consequently, this rule delivers a reporting guarantee, ensuring with high confidence that the actual welfare delivered exceeds a lower threshold, while also having the favorable regret properties of all risk-aware rules. The proposed rule thus ensures Policy Learning with Confidence, and is subsequently referred to as the PoLeCe rule.
@@
To illustrate, consider a setting in which a budget-conscious DM needs to allocate funds across several social programs that affect the same group of individuals, either in response to a slight budget increase or for incremental cost-cutting. Suppose the DM wishes to use the Marginal Value of Public Funds (MVPF), spearheaded by Hendren and Sprung-Keyser (2020), for this purpose. The MVPF of each social program is the ratio of its marginal benefit to its net marginal cost to the government, inclusive of the impact of any behavioral responses on the government budget. For instance, an MVPF of 1.5 indicates that each \$1 of net government spending generates \$1.50 in benefits for beneficiaries. For a utilitarian DM, the MVPF provides a metric to compare the “bang for the buck” of different policies that target the same beneficiaries. A DM interested in allocating additional marginal expenditure will prioritize programs with the highest MVPF to maximize welfare. [^3] Here we abstract from distributional incidence and assume the DM values dollars equally across the beneficiaries of each policy. If welfare weights vary across beneficiaries, as in the original framework of Hendren and Sprung-Keyser (2020), our analysis can accommodate this by adjusting each MVPF in the value function using the DM’s designated weights. Example welfare weights have been derived in Hendren (2020). Other metrics for measuring programs’ marginal benefits and costs have also been discussed in García and Heckman (2022).